# Anomalous strain rate sensitivity of a Cu/Al2O3 multi-layered thin film

### v1

@@ Abstract
To study the size and strain rate dependency of copper polycrystalline microstructures, a multi-layered copper/Al2O3 thin film was deposited on a Si substrate using a hybrid deposition system (combining physical vapour and atomic layer deposition). High temperature treatment was applied on the “As Deposited” material with ultrafine-grained structure to increase the average grain size, resulting in a “Heat Treated” state with microcrystalline structure. Focused ion beam milling was employed to create square shaped micropillars with two different sizes, that were subjected to compressive loading at various (0.001/s – 1000/s) strain rates. The detected two distinct anomalies in the strain rate sensitivity behavior appearing at high strain rates could be related to the pillar diameter and the grain size of the deformed samples. The Al2O3 interlayer studied by transmission electron microscopy showed excellent thermal stability and grain boundary pinning by precipitation, also resulting in the homogeneous deformation of the pillars and preventing shear localization. Geometrically necessary dislocation densities estimated by high (angular) resolution electron backscatter diffraction presented inhomogeneous dislocation distribution within the deformed pillar volumes, that is attributed to the proximity of the sample edges. Finally, the Al2O3 interlayers successfully suppressed any possible recrystallization processes, contributing to the excellent film stability, that makes the proposed coating ideal to be operating under extreme conditions.
@@
journal: ArXiv.org
@@ 1 Introduction
The aim of this study is, therefore, to design a hybrid (Cu/Al2O3) thin film and tailor its microstructure by heat treatment to become a model material in order to study the SRS behavior as a function of sample size and microstructure. The purpose of the Al2O3 interlayer was to limit grain growth between the subsequent Cu depositions, allowing to investigate the “As Deposited” (ultrafine-grained) coating parallel to its “Heat Treated” counterpart containing microcrystalline grains.
@@ 2.1 Sample preparation
For the creation of the desired multilayers, a hybrid thin film deposition system combining both ALD and PVD methods was applied (Swiss Cluster). The sequential deposition consisted of alternating (5 times) $\sim 1$ $\upmu$m thick pure Cu separated by 10 nm thin Al2O3 interlayers deposited on a Si (100) substrate (Fig. 1a) at room temperature. Cross-sectional electron backscatter diffraction (EBSD) confirmed the average grain size of this “As Deposited - AD” film to be $\overline{d^{AD}}\sim 120$ nm (area weighted mean, see in Fig. 1c,e), qualifying this coating as ultrafine-grained (UFG).
@@ 2.2 Micromechanical testing
Micro-samples were deformed using an Alemnis ASA nanodeformation system equipped with three different load cells (strain gage based standard and mini load cells for quasi-static strain rates (SR), piezo-based SmarTip for high SRs). For pillar compressions, conductive diamond flat punches (Synton AG) with diameters of 5 $\upmu$m and 10 $\upmu$m were utilized. The recorded load-displacement curves were baseline (drift) and compliance corrected using the AMMDA evaluation software. During HSR testing, due to the high acceleration following the tip actuation (achieving up to 6 mm/s tip velocity), an oscillating noise appeared, superimposed to the signal (“ringing effect”). In order to determine the yield stresses $\sigma_{\mathrm{y}}$ in such difficult cases, the elastic and plastic parts of the stress-strain curves were fitted using a linear equation, and their intersection was used to obtain the apparent yield stress (see Fig. 2a,b and Suppl. Fig. S4 for the full process). This method enabled to avoid further data filtering or averaging in the $\sigma_{\mathrm{y}}$ determination. For the consistency of the yield stress values at low and high strain rates, all mechanical data were treated the same way.
@@ 3 Results
Smaller pillars exhibit lower yielding, that can be attributed to the i) surface damage caused by FIB milling [21, 22] or ii) it can be a grain size effect enhancing strain localization in the smaller HT specimen. As it was mentioned previously, to evade tapering, a higher incident angle of the bombarding ions were utilized, that may result in a higher concentration of implanted Ga close to the pillar surfaces. In smaller pillars, the effect of Ga presence is enhanced due to the increased surface-to-volume ratio. This explanation is the most probable cause of the slight yield stress decrease in smaller pillars rather that other microstructure related effects, since it is apparent in both the AD (UFG) and HT (MC) pillars in a similar manner at both low and high strain rates (see in Fig. 3).
@@
In Fig. 3a, the determined average $\sigma_{\mathrm{y}}$ values are plotted on a linear scale as a function of $\dot{\varepsilon}$ on a logarithmic scale. The error bars show the spreading of the calculated values from repeated pillar compression tests (typically three), that presents good repeatability of the experiments at almost all applied strain rates. As a result of the employed linear fitting method, $\sigma_{\mathrm{y}}$ can be determined fairly consistently. The shaded area represents the noise during loading, indicating that when a load cell is switched from strain gage based ($\dot{\varepsilon}\leq$ 1/s) to piezoelectric based detection ($\dot{\varepsilon}\geq$ 10/s), the noise increases significantly. The drawback of using a noisier (but much stiffer) load cell [10] is compensated by the ability to go beyond the previously achievable deformation speeds and sampling rates. The noise becomes extremely large then the load signal starts to have a overwhelming oscillation due to the ringing effect ($\dot{\varepsilon}=$ 1000/s). To reduce this outcome originating from the immense acceleration during the operation (6 mm/s), certain pillars were compressed in a way that the tip acceleration was initiated far ($\sim 14$ $\upmu$m) away from the surface, anticipating that at this distance the tip resonance decreases to minimal upon impact, when the flat punch is already travelling at its maximum speed [23]. To further improve the mechanical data, some pillars were compressed to very high strains to allow the tip to drive through these pillars at a constant speed without stopping, thereby completely destroying the samples and preventing them from being examined by post mortem EBSD analysis (hence the notation “pancake” in Suppl. Fig. S5-S7).
@@ 4 Discussion
Slip plane localization is successfully suppressed by the consecutive Al2O3 layers in both the AD and HT samples, hence the resulting smooth $\sigma_{\mathrm{eng}}$ $\varepsilon_{\mathrm{eng}}$ curves in Fig. 2, that show the absence of large strain bursts (or stress drops). The surface analysis of the deformed pillars support these findings in Suppl. Figs. S2 and S3. Flow stresses reach quite high average values ($\overline{\sigma_{\mathrm{y}}^{\mathrm{AD}}}=758$ MPa and $\overline{\sigma_{\mathrm{y}}^{\mathrm{HT}}}=460$ MPa) during quasi-static deformation, much higher than pure Cu in literature, approaching the regime of nanocrystalline copper [24, 25]. This increased yielding could be partly attributed to the grain boundary (Hall-Petch) strengthening (using the already established $\overline{d}$ values for the AD and HT samples with the equation $\sigma_{\mathrm{GB}}=\sigma_{0}+kd^{-0.5}$ with $\sigma_{0}^{\mathrm{Cu}}=25$ MPa material constant and $k=0.12$ MPa m1/2 strengthening coefficient [26], resulting in $\sigma_{\mathrm{GB}}^{\mathrm{AD}}=371$ MPa and $\sigma_{\mathrm{GB}}^{\mathrm{HT}}=139$ MPa theoretical values). Further strengthening mechanisms also play a key role in raising the $\sigma_{\mathrm{y}}$ values, such as dislocation and potentially (based on Fig. 4c) precipitation strengthening. As suggested by Fig. 4a, nanotwinning also occurs in this thin film system [27] that could be significant in the HSR behavior in case of possible dynamic recrystallization [28]. Additionally, high angle grain boundaries in the AD sample tend to run along the deformation axis of the pillars (somewhat columnar grain growth resulting in ellipsoid grain shapes), that will also increase the material’s resistance towards compressive/tensile stresses.
@@
When we compare the yield stress values to the 3D printed Cu UFG/MC system published by Ramachandramoorthy and Kalacska et al. ([18]), we found these results to be in good agreement with both the AD/HT yielding, meaning that the multi-layered thin film proposed here is a fitting model material to further elaborate on the anomalous behavior of the polycrystalline Cu strain rate sensitivity.
@@
In Fig. 3, two anomalous trends in the SRS can be observed: the first one being the saturation of $\sigma_{\mathrm{y}}$ values of the AD samples at HSR in case of the smaller pillars. This behavior was already reported in Ref. [18], where circular Cu polycrystalline pillars with 2 $\upmu$m gauge diameters were tested. There it was found that only the pillars with UFG structure exhibited such anomaly, modifying the SRS factor from $m_{\mathrm{UFG}}^{\mathrm{LSR}}=0.06$ to $m_{\mathrm{UFG}}^{\mathrm{HSR}}=0.003$. Here, we detected similar behavior, but remarkably, only in the smaller (2.9 $\upmu$m side length) pillars, leading to the conclusion that the pillar microstructure is not the only driving factor in the anomalous SRS behavior, but also the size of the sample has an additional influence too. In our AD thin films, the SRS factors were found to be close to the already reported values ($m_{\mathrm{AD}}^{\mathrm{LSR}}=0.03$, $m_{\mathrm{AD}}^{\mathrm{HSR}}=0.01$). When pillars were fabricated with bigger (4 $\upmu$m) side length, this anomalous behavior disappeared, and a single $m$ value could describe the SRS within the whole investigated $\dot{\varepsilon}$ regime.
The apparent activation volume at the yield point is calculated by the following equation [29, 30]:
@@
where $k_{\mathrm{B}}$ and $T$ are the Boltzmann constant and the temperature, respectively. By using $b=2.56\mathrm{\AA}$ as the Burgers vector of a {111}$<110>$ dislocation in copper, the calculated activation volumes are $\Omega_{\mathrm{AD}}^{\mathrm{big}}\sim 7.5b^{3}$, $\Omega_{\mathrm{HT}}^{\mathrm{LSR}}\sim 5.5b^{3}$ and $\Omega_{\mathrm{AD}}^{\mathrm{small-HSR}}\sim 13b^{3}$. This suggests that in the LSR regime, the deformation is controlled by the movement of dislocations ($1b^{3}-10b^{3}$) nucleated at high angle grain boundaries and at surface sources. In the small AD pillars, the increase in $\Omega$ suggests a collective dislocation dynamic process ($10b^{3}-100b^{3}$) [31], meaning that dislocations are created at multiple sources in large numbers upon impact, but their movement is restrained.
@@
The second anomaly in Fig. 3 was discovered upon testing the pillars above 100/s, where the $\sigma_{\mathrm{y}}$ values unexpectedly increase in the HT samples, regardless of their size. At LSRs, HT pillars exhibited similar SRS ($m_{\mathrm{HT}}^{\mathrm{LSR}}=0.02$) as the AD samples. This value matches the $m$ factor reported in Ref. [18] for the MC structured pillars, where a monotonous SRS was also observed in this $\dot{\varepsilon}$ regime. However, upon testing the HT pillars at $\dot{\varepsilon}=1000$/s, the detected $m$ value has increased significantly ($m_{\mathrm{HT}}^{\mathrm{big-HSR}}=0.19$, $m_{\mathrm{HT}}^{\mathrm{small-HSR}}=0.11$), showing that at HSR, a transition in the dislocation dynamics happen, resulting in a strain rate hardening regime. This crossover has already been proposed by simulations [32] and demonstrated by large scale impact testing [33], but, to our knowledge, it is the first time to be presented at such small scale, during micromechanical testing. According to Ref. [33], at LSRs, dislocations are thermally activated, while when $\dot{\varepsilon}$ exceeds $10^{2}$/s, dislocation velocities become so high that the phonon-drag (thermal elastic vibrations) on their motion can no longer be neglected [34], making dislocations much more difficult to move, due to which we observe such an outstanding increase in the $\sigma_{\mathrm{y}}$ values. Just like in case of the first anomaly, the microstructure plays a crucial part in the SRS, as this deviation will most likely present itself in the AD samples too, only at somewhat higher SRs than what is currently reachable by small scale experiments [35].

### v2

# Strain rate sensitivity of a Cu/Al2O3 multi-layered thin film

@@ Abstract
To study the size and strain rate dependency of copper polycrystalline microstructures, a multi-layered copper/Al2O3 thin film was deposited on a Si substrate using a hybrid deposition system (combining physical vapour and atomic layer deposition). High temperature treatment was applied on the “As Deposited” material with ultrafine-grained structure to increase the average grain size, resulting in a “Heat Treated” state with microcrystalline structure. Focused ion beam milling was employed to create square shaped micropillars with two different sizes, that were subjected to compressive loading at various (0.001/s – 1000/s) strain rates. Differences in the strain rate sensitivity behavior manifesting at low and high strain rates are discussed in the context of the pillar diameters and the grain size of the deformed samples. The Al2O3 interlayer studied by transmission electron microscopy showed excellent thermal stability and grain boundary pinning by precipitation, also resulting in the homogeneous deformation of the pillars and preventing shear localization. Geometrically necessary dislocation densities estimated by high (angular) resolution electron backscatter diffraction presented inhomogeneous dislocation distribution within the deformed pillar volumes, that is attributed to the proximity of the sample edges. Finally, the Al2O3 interlayers successfully suppressed any possible recrystallization processes, contributing to the excellent film stability, that makes the proposed coating ideal to be operating under extreme conditions.
@@
journal: Surface and Coatings Technology
@@ 1 Introduction
The aim of this study is, therefore, to design a hybrid (Cu/Al2O3) thin film and tailor its microstructure by heat treatment to study the SRS behavior as a function of sample size and microstructure. The applied material can be considered as “model material”, since the aim of this work was to provide a straightforward method for studying certain high strain rate properties (e.g. strain rate sensitivity of the yield stress as a function of microstructure), due to Cu’s well-understood characteristics. Since the preparation methodology can be consistently repeated, the fundamental properties and underlying mechanisms can be representative to a broader class of thin films, making it a useful proxy for understanding similar systems. The purpose of the Al2O3 interlayer was to limit grain growth between the subsequent Cu depositions, allowing to investigate the “As Deposited” (ultrafine-grained) coating parallel to its “Heat Treated” counterpart containing microcrystalline grains.
@@ 2.1 Sample preparation
For the creation of the desired multilayers, a hybrid thin film deposition system combining both ALD and PVD methods was applied (Swiss Cluster). The sequential deposition consisted of alternating (5 times) $\sim 1$ $\upmu$m thick pure Cu separated by 10 nm thin Al2O3 interlayers, deposited on a Si (100) substrate coated with a 100 nm thick SiO2 diffusion barrier (Fig. 1a) at room temperature. Copper was deposited by DC magnetron sputtering in current control mode. Two magnetrons were used, identical by design, mounted 180 apart within the chamber. The current on both was set to 200 mA, which resulted in 350 V of plasma potential and an average of 70 W plasma power. The base pressure prior deposition was better than $5\times 10^{-7}$ mbar. Targets with 50 mm diameter and 3 mm thickness were sputtered, supplied by HMW-Hauner GmbH, with a purity of 99.99%. The substrate was kept at room temperature. During deposition, 15 sccm argon was released into the chamber, and the process pressure was held at $5\times 10^{-3}$ mbar by a throttle valve. It is important to note that the key advantage of the Swiss Cluster equipment is that the vacuum was maintained between PVD and ALD transitions, preventing the formation of a native oxide layer. Furthermore, these layers were prepared using the classical ALD method by alternating trimethylaluminum (TMA) and water as precursors [19]. This methodology was employed deliberately as a baseline for comparison and validation, adhering to widely accepted practices in the field [20, 21]. Cross-sectional electron backscatter diffraction (EBSD) confirmed the average grain size of this “As Deposited - AD” film to be $\overline{d^{AD}}\sim 120$ nm (area weighted mean, see in Fig. 1c,e), qualifying this coating as ultrafine-grained (UFG).
@@ 2.2 Micromechanical testing
Micro-samples were deformed using an Alemnis ASA nanodeformation system equipped with three different load cells (strain gage based standard and mini load cells for quasi-static strain rates (SR), piezo-based SmarTip for high SRs). For pillar compressions, conductive diamond flat punches (Synton AG) with diameters of 5 $\upmu$m and 10 $\upmu$m were utilized. The recorded load-displacement curves were baseline (drift) and compliance corrected using the AMMDA evaluation software. During HSR testing, due to the high acceleration following the tip actuation (achieving up to 6 mm/s tip velocity), an oscillating noise appeared, superimposed to the signal (“ringing effect”). In order to determine the yield stresses $\sigma_{\mathrm{y}}$ in such difficult cases, the elastic and plastic parts of the stress-strain curves were fitted using a linear equation, and their intersection was used to obtain the apparent yield stress (see Fig. 2a,b and Suppl. Fig. S4 for the full process). This method enabled to avoid further data filtering or averaging in the $\sigma_{\mathrm{y}}$ determination. For the consistency of the yield stress values at low and high strain rates, all mechanical data were treated the same way. Additionally, the more common 1% plastic strain offset method was also employed to assess the SRS behaviour, that is provided in Suppl. Fig. S5 for comparison.
@@ 3 Results
Smaller pillars exhibit lower yielding, that can be attributed to the i) surface damage caused by FIB milling [24, 25] or ii) it can be a grain size effect enhancing strain localization in the smaller HT specimen. As it was mentioned previously, to evade tapering, a higher incident angle of the bombarding ions were utilized, that may result in a higher concentration of implanted Ga close to the pillar surfaces. In smaller pillars, the effect of Ga presence is enhanced due to the increased surface-to-volume ratio. This explanation is the most probable cause of the slight yield stress decrease in smaller pillars rather than other microstructure related effects, since it is apparent in both the AD (UFG) and HT (MC) pillars in a similar manner at both low and high strain rates (see in Fig. 3).
@@
In Fig. 3a, the determined average $\sigma_{\mathrm{y}}$ values are plotted on a linear scale as a function of $\dot{\varepsilon}$ on a logarithmic scale. The error bars show the spreading of the calculated values from repeated pillar compression tests (typically three), that presents good repeatability of the experiments at almost all applied strain rates. As a result of the employed linear fitting method, $\sigma_{\mathrm{y}}$ can be determined fairly consistently. The shaded area represents the noise during loading, indicating that when a load cell is switched from strain gage based ($\dot{\varepsilon}\leq$ 1/s) to piezoelectric based detection ($\dot{\varepsilon}\geq$ 10/s), the noise increases significantly. The drawback of using a noisier (but much stiffer) load cell [10] is compensated by the ability to go beyond the previously achievable deformation speeds and sampling rates. The noise becomes extremely large then the load signal starts to have a overwhelming oscillation due to the ringing effect ($\dot{\varepsilon}=$ 1000/s). To reduce this outcome originating from the immense acceleration during the operation (reaching 6 mm/s tip velocity upon impact), certain pillars were compressed in a way that the tip acceleration was initiated far ($\sim 14$ $\upmu$m) away from the surface, anticipating that at this distance the tip resonance decreases to minimal upon impact, when the flat punch is already travelling at its maximum speed [26]. To further improve the mechanical data, some pillars were compressed to very high strains to allow the tip to drive through these pillars at a constant speed without stopping, thereby completely destroying the samples and preventing them from being examined by post mortem EBSD analysis (hence the notation “pancake” in Suppl. Fig. S6-S8).
@@ 4 Discussion
Slip plane localization is successfully suppressed by the consecutive Al2O3 layers in both the AD and HT samples, hence the resulting smooth $\sigma_{\mathrm{eng}}$$\varepsilon_{\mathrm{eng}}$ curves in Fig. 2, that show the absence of large strain bursts (or stress drops). The surface analysis of the deformed pillars support these findings in Suppl. Figs. S2 and S3. Flow stresses reach quite high average values ($\overline{\sigma_{\mathrm{y}}^{\mathrm{AD}}}=758$ MPa and $\overline{\sigma_{\mathrm{y}}^{\mathrm{HT}}}=460$ MPa) during quasi-static deformation, much higher than pure Cu in literature, approaching the regime of nanocrystalline copper [27, 28]. This increased yielding could be partly attributed to the grain boundary (Hall-Petch) strengthening (using the already established $\overline{d}$ values for the AD and HT samples with the equation $\sigma_{\mathrm{GB}}=\sigma_{0}+kd^{-0.5}$, with $\sigma_{0}^{\mathrm{Cu}}=25$ MPa material constant and $k=0.12$ MPa m1/2 strengthening coefficient [29], resulting in $\sigma_{\mathrm{GB}}^{\mathrm{AD}}=371$ MPa and $\sigma_{\mathrm{GB}}^{\mathrm{HT}}=139$ MPa theoretical values). Further strengthening mechanisms also play a key role in raising the $\sigma_{\mathrm{y}}$ values, such as dislocation and potentially (based on Fig. 4c) precipitation strengthening. As suggested by Fig. 4a, nanotwinning also occurs in this thin film system [30] that could be significant in the HSR behavior in case of possible dynamic recrystallization [31]. Additionally, high angle grain boundaries in the AD sample tend to run along the deformation axis of the pillars (somewhat columnar grain growth resulting in ellipsoid grain shapes), that will also increase the material’s resistance towards compressive/tensile stresses.
@@
When we compare the yield stress values to the 3D printed Cu UFG/MC system published in Ref. [18], we found these results to be in good agreement with both the AD/HT yielding, meaning that the multi-layered thin film proposed here is a fitting model material to further elaborate on the previously detected anomalous behavior of the polycrystalline Cu strain rate sensitivity.
@@
In Fig. 3, two trends in the SRS is discussed: i) the monotonous $\sigma_{\mathrm{y}}$ increase with increasing SR in all studied samples at lower strain rates, and ii) $\sigma_{\mathrm{y}}$ increase at 1000/s. The first behavior was already reported in the literature on similar materials (e.g. [18], where circular Cu polycrystalline pillars with 2 $\upmu$m gauge diameters were tested). There, it was found that pillars with UFG structure exhibited an anomalous SR saturation at $\dot{\varepsilon}\sim$ 500/s, modifying the SRS factor from $m_{\mathrm{UFG}}^{\mathrm{LSR}}=0.06$ to $m_{\mathrm{UFG}}^{\mathrm{HSR}}=0.003$. Here, we found similar behavior looking at the AD results between $0.1$ $s^{-1}$$<\dot{\varepsilon}\leq 100$ $s^{-1}$ (Fig. 3a). However, $m$ remained constant in the smaller (2.9 $\upmu$m side length) pillars at HSR, leading to the conclusion that the pillar microstructure is not the only driving factor in the SRS behavior, but also the size of the sample has an additional influence too. In our AD thin films, the SRS factors were found to be close to the already reported values ($m_{\mathrm{AD}}^{\mathrm{LSR}}=0.03$, $m_{\mathrm{AD}}^{\mathrm{HSR}}=0.01$). When pillars were fabricated with bigger (4 $\upmu$m) side length, the SRS factor $m$ remained constant throughout the whole investigated $\dot{\varepsilon}$ regime.
Assuming that the active deformation mechanisms are temperature activated allows one to determine the activation volume characterizing the stress dependent part of the Gibbs potential. If deformation is governed by a single mechanism then the activation volume is expected to be constant at different applied stresses. The change in the activation volume, therefore, suggests a change in the dominant deformation mechanism. The apparent activation volume at the yield point is calculated by the following equation [32, 33, 34]:
@@
where $k_{\mathrm{B}}$ and $T$ are the Boltzmann constant and the temperature, respectively. By using $b=0.256$ nm as the Burgers vector of a {111}$<110>$ dislocation in copper, and $T=297$ K, the calculated activation volumes are $\Omega_{\mathrm{AD}}^{\mathrm{big}}\sim 15b^{3}$, $\Omega_{\mathrm{AD}}^{\mathrm{small}}\sim 49b^{3}$, and $\Omega_{\mathrm{HT}}^{\mathrm{small-HSR}}\sim 8b^{3}$. The measured volumes are in the same range as previous measurements on similar (UFG) materials [14, 35]. Since in materials with complex microstructure, like the one studied here, usually several deformation mechanisms are active simultaneously, interpretation of activation volumes is rather challenging. If one assumes that there is a dominant mechanism, it suggests that in the LSR regime in the HT pillars, the deformation is controlled by the movement of dislocations ($1b^{3}-10b^{3}$) nucleated at high angle grain boundaries and at surface sources [36]. In the small AD pillars, the increase in $\Omega$ suggests a collective dislocation dynamic process ($10b^{3}-100b^{3}$) [36], meaning that dislocations are created at multiple sources in large numbers upon impact, but their movement is restrained. To draw more detailed conclusions one would need to gain more information about the active deformation mechanisms either by, e.g., performing an in situ TEM experiment [34] or by analyzing complementary atomistic simulations [37]. Other mechanisms could also play a significant role in the SRS behaviour (e.g. grain boundary diffusion-assisted processes influenced by creep-deformation [38, 39]), that requires more targeted experimental studies to be performed.
@@
Looking at Fig. 3it can be seen that upon testing the pillars above 100/s, where the $\sigma_{\mathrm{y}}$ values unexpectedly increase in the HT samples, regardless of their size. Firstly, this apparent SRS change can originate from the increased noise levels during HSR testing. At LSRs, HT pillars exhibited similar SRS ($m_{\mathrm{HT}}^{\mathrm{LSR}}=0.02$) as the AD samples. This value matches the $m$ factor reported in Ref. [18] for the MC structured pillars, where a monotonous SRS was also observed in this $\dot{\varepsilon}$ regime. However, upon testing the HT pillars at $\dot{\varepsilon}=1000$/s, the detected $m$ value has increased significantly ($m_{\mathrm{HT}}^{\mathrm{big-HSR}}=0.19$, $m_{\mathrm{HT}}^{\mathrm{small-HSR}}=0.11$), showing that at HSR, a transition in the dislocation dynamics may happen, resulting in a strain rate hardening regime. This crossover has already been proposed by simulations [40] and demonstrated by large scale impact testing [41, 42], but, to our knowledge, it is the first time to be presented at such small scale, during micromechanical testing. According to Ref. [42], at LSRs, dislocations are thermally activated, while when $\dot{\varepsilon}$ exceeds $10^{2}$/s, dislocation velocities become so high that the phonon-drag (thermal elastic vibrations) on their motion can no longer be neglected [43], making dislocations much more difficult to move, due to which we observe such an outstanding increase in the $\sigma_{\mathrm{y}}$ values. As the microstructure (and the initial dislocation density) plays a crucial part in the SRS, if such a mechanism is anticipated, this deviation will most likely present itself in the AD samples too, only at somewhat higher SRs than what is currently reachable by small scale experiments [44]. However, to have a definitive conclusion of the origin of this phenomenon, a more rigorous and systematic HSR testing is required.